Provide formulas and procedure in excel

QUESTION 9

  1. Marst Corporation's budgeted production in units and budgeted raw materials purchases over the next three months are given below:

January

February

March

Budgeted production (in units)

70,200

????

81,000

Budgeted raw materials (in pounds)

143,400

154,600

159,800

  1. Four pounds of raw materials are required to produce one unit of product. The company wants raw materials on hand at the end of each month equal to 31% of the following month's production needs. The company is expected to have 42,400 pounds of raw materials on hand on January 1.
  2. Required:
  3. Prepare the Budgeted production for February
Answer #1

ANSWER:

ANSWER IS 19,637.68 UNITS

Working Notes:
Let X be the Budgeted production units.
So, Beginning raw material Inventory = February budgeted production units x 4 x 31%
= X x 4x 31%
=1.24X
Ending Raw material inventory = March budgeted production units x 4 x 31%
=81,000 x 4 x 31%
=100,440
February production required raw material = Budgeted production units. x 4
= 4X
Budgeted raw materials purchases for February = Ending Raw material inventory + February production required raw material - Beginning raw material Inventory
154,600 = 100,440 + 4X - 1.24X
154,600= 100,400 + 2.76X
X=(154,600 - 100,400)/2.76
X=54,200/2.76
X=19,637.68 units

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU*************